Please help me because I am really struggling with this concept! Thank you!
4/5y + y-2/y = -1/5

Answer by Alan3354(69443)

Please help me because I am really struggling with this concept! Thank you!
4/5y + y-2/y = -1/5
-------------------
If you mean (y-2)/y for the 2nd term:
4/5y + y-2/y = -1/5
Multiply thru by 5y
4 + 5(y-2) = -y
4 + 5y-10 = -y
4 + 6y = 10
6y = 6
y = 1